The sum of a boy and his father's age is 63 years. The difference between their ages is 33 years....Translate the two statements and solve for each variable.
Answer by vidhyak(98)

Let boy's age be b
Let father's age be f

The sum of a boy and his father's age is 63 years
f + b = 63 -----------------------eq. 1

The difference between their ages is 33 years
f - b = 33 -----------------------eq. 2


Solving the 2 equations(Add the 2 equations)
f + b = 63
f - b = 33

2f = 96
f = 96/2 = 48

Father's age is 48
Substitute f = 48 in eq 1

f + b = 63
48 + b = 63
b = 63 - 48
b = 15

Boy's age is 15
Father's age is 48